# onlineReferences

### v5

@@ Appendix C Decomposing the TWFEIV estimand in Oreopoulos (2006)
[…]
@@
and $\hat{Z}_{e,t}$ is the residuals from regressing $Z_{i,t}$ on the cohort and North Ireland fixed effects [^24] We can write $\hat{Z}_{i,t}=\hat{Z}_{e,t}$ because $Z_{i,t}$ only varies across cohorts and regions..
[…]
$\beta_{IV}=$ $\sum\limits_{U-1\geq t\geq e}w_{e,t}^{1}\cdot CLATT_{e,t}+\sum\limits_{t\geq U}w_{e,t}^{2}\cdot\Delta_{e,t},$ (20)
@@
$\Delta_{e,t}=\displaystyle\frac{CAET^{1}_{e,t}\cdot CLATT_{e,t}-CAET^{1}_{U,t}\cdot CLATT_{U,t}}{CAET^{1}_{e,t}-CAET^{1}_{U,t}},$
@@
$=\frac{E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot(CAET^{1}_{e,t}-CATT^{1}_{U,t})}{\sum\limits_{U-1\geq t\geq e}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot CAET^{1}_{e,t}+\sum\limits_{t\geq U}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot(CAET^{1}_{e,t}-CAET^{1}_{U,t})},$ (24)
@@
where $n_{e,t}$ and $E[\hat{Z}_{i,t}|E_{i}=e]$ represent the population share and residuals for England in cohort $t$ respectively.
[…]
@@
Figure 3 plots the weight and the corresponding estimate for each $CLATT_{e,t}$ and $\Delta_{e,t}$ (bias term). The consistent estimators for each $w^{1}_{e,t}$, $w^{2}_{e,t}$, $CLATT_{e,t}$ and $\Delta_{e,t}$ are constructed from the sample analogue, exploiting $\widehat{WDID}_{e,t}$ and $(D_{e,t}-D_{e,1946})-(D_{U,t}-D_{U,1946})$.
[…]
@@
Figure 3: Decomposition result for the TWFEIV estimand in Oreopoulos (2006). Notes: This figure plots the estimated weights and the corresponding estimates for each $CLATT_{e,t}$ and $\Delta_{e,t}$ (bias term). The closed squares are the weights and the corresponding estimates for each $CLATT_{e,t}$. The closed triangles are the weights and the corresponding estimates for each $\Delta_{e,t}$ (bias term).
Appendix D Relationship between DID and DID-IV
In this appendix, we describe the relationship between DID and DID-IV designs, and establish that DID-IV is DID with noncompliance of the assigned treatment.
Following Subsection 2.1, we introduce the notation in DID designs. Let $Z_{i}=(Z_{i,0},Z_{i,1})$ denote the path of a program that distributes the treatment $D_{i,t}$, and let $D_{i}=(D_{i,0},D_{i,1})$ denote the path of the treatment for unit $i$. In the canonical DID set up, $Z_{i}=(Z_{i,0},Z_{i,1})$ takes only two values: $Z_{i}=(0,1)$ if unit $i$ is assigned to the program in period $1$ (treatment group) and $Z_{i}=(0,0)$ if unit $i$ is not assigned to the program in period $1$ (control group). Let $E_{i}\in\{0,1\}$ denote the group variable: $E_{i}=1$ if unit $i$ belongs to the treatment group and $E_{i}=0$ if unit $i$ belongs to the control group. We refer to treatment and control groups as exposed and unexposed groups, respectively.
Note that in the common DID set up, the treatment status $D_{i,t}$ is equal to the exposure to the program $Z_{i,t}$ because we ex ante impose two restrictions on the treatment adoption behavior across units. First, we assume that all units are not allowed to receive the treatment in period 00.
$D_{i,0}((0,1))=D_{i,0}((0,0))=0\hskip 8.53581pt\text{for all $i$}.$ (25)
Second, we assume that unit $i$ should receive the treatment in period $1$ if that unit belongs to an exposed group, and that it should not receive the treatment in period $1$ if that unit belongs to an unexposed group.
$1=D_{i,1}((0,1))>D_{i,1}((0,0))=0\hskip 8.53581pt\text{for all $i$}.$ (26)
These restrictions imply that we have Tables 5-6 under DID designs: there exists only the type $CM^{Z}\land NT^{T}$ in exposed and unexposed groups, respectively.
Then, under these restrictions, we can show that DID-IV designs nest DID designs as a special case along with three dimensions. First, Assumptions 3-5 in DID-IV designs are automatically satisfied in DID designs. Assumption 3 follows from (26) and Assumption 4 follows from (25). Assumption 5 follows from combining (25) with (26).
Second, the target parameter in DID designs, the ATT is equal to the LATET in the canonical DID set up; this follows from Tables 5-6:
Table 5: Exposed group (assigned to $z=(0,1)$)
observed counterfactual
$(D_{0}((0,0)),D_{1}((0,1)))$ $D_{1}((0,0))=0$
$D_{0}((0,0))=0,D_{1}((0,1))=1$ $CM^{Z}\land NT^{T}$
Table 6: Treatment group (assigned to $z=(0,0)$)
observed counterfactual
$(D_{0}((0,0)),D_{1}((0,0)))$ $D_{1}((0,1))=1$
$D_{0}((0,0))=0,D_{1}((0,0))=0$ $CM^{Z}\land NT^{T}$
$ATT$ $=E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1]$
$=E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1,CM^{Z}]$
$=LATET.$
Note that both ATT and LATET measure the treatment effects, for those who are in an exposed group (assigned to $z=(0,1)$) and are induced to treatment by instrument ($CM^{Z}$). Therefore, we define the LATET as our target parameter in DID-IV designs in Subsection 2.2.
Finally, a well-known parallel trends assumption in DID designs can also be interpreted as requiring the same expected time gain between exposed and unexposed groups. From Tables 5-6, we can rewrite the parallel trends assumption in the untreated outcome as:
$E[Y_{i,1}(0)-Y_{i,0}(0)|E_{i}=0]=E[Y_{i,1}(0)-Y_{i,0}(0)|E_{i}=1]$
$\iff$ $E[Y_{i,1}(0)-Y_{i,0}(0)|E_{i}=0,NT^{T}]=E[Y_{i,1}(0)-Y_{i,0}(0)|E_{i}=1,NT^{T}]$
$\iff$ $\Delta_{NT^{T},0}=\Delta_{NT^{T},1}$
$\iff$ $\Delta_{0}=\Delta_{1},$
where the last equivalence holds because we have only the time never takers $NT^{T}$ in exposed and unexposed groups respectively under DID designs.
$w_{NT^{T},0}=Pr(NT^{T}|E_{i}=e)=1,\hskip 8.53581pt\text{for all}\hskip 8.53581pte=0,1.$
[…]
Appendix E Comparing DID-IV to Fuzzy DID
[…]
E.1   Comparing the identifying assumptions in dCDH and this article
In this subsection, we investigate the detailed connections between dCDH’s identifying assumptions and ours. Because dCDH consider the repeated cross section settings, when we compare the identifying assumptions between the two papers, we refer to Assumptions 14-19. For now, we ignore the difference in the definition of target parameter.
dCDH consider the following identification assumptions for the Wald-DID estimand to capture the SLATET. [^25] In repeated cross section settings, the formal definition is provided in Assumption 31. Note that in dCDH, a group variable $G$ plays the role of instrument $Z$ in our notation. [^26] This follows from two observations in dCDH. First, dCDH call the groups $G=1$ and $G=0$ treatment and control groups, respectively. Second, the group variable $G$ is included in Assumption 28 (Treatment participation equation). We therefore replace $G$ with $Z$ and call the groups $Z=1$ and $Z=0$ exposed and unexposed groups, respectively. [^27] Note that in Appendix A.2, we distinguish the group variable $E$ from the instrument path $Z$ to uncover Assumption 16 (No anticipation in the first stage). For the detailed explanations for each assumption, see dCDH.
@@ E.1   Comparing the identifying assumptions in dCDH and this article
[…]
@@
In Assumption 30, we modify the dCDH’s notation $D(t)$ to ease its interpretation: we define $D_{t}(Z)=\mathbf{1}\{V\geq v_{Zt}\}$ to be the treatment status under time $T=t$ in group $Z$ instead of defining it as $D(t)$.
[…]
Finally, dCDH implicitly impose Assumption 14 and Assumption 16 when they formalize potential outcomes and treatment participation equations. Specifically, dCDH do not introduce the instrument path in their framework and define potential outcomes as $Y(D)$, which implies Assumption 14. As one can see in Assumption 28, dCDH assume that the treatment participation equation only depends on the current instrument status, which implies Assumption 16 since dCDH assume the sharp assignment of the instrument: $v_{10}=v_{00}$.
The differences between dCDH’s assumptions and ours are as follows. First, dCDH assume the parallel trends assumption in the untreated outcome (Assumption 29), whereas we assume the parallel trends assumption in the unexposed outcome. Second, dCDH assume the stable treatment rate assumption in an unexposed group (Assumption 27), whereas we assume the parallel trends assumption in the treatment between exposed and unexposed groups. Third, dCDH impose the independence assumption between unobserved heterogeneity $V$ and time variable $T$ given $Z$ in Assumption 28, whereas we do not impose such restriction. Forth, dCDH argue that the Wald-DID estimand requires the stable treatment effect assumption during the two periods (Assumption 30) to capture the SLATET, whereas we do not require such restriction for this estimand to identify the LATET. Finally, dCDH assume the condition (27) in Assumption 26, which requires that the treatment rate in an exposed group increases between the two periods, whereas we do not assume this condition.
Remark E.1.1.
@@
where we define282828dCDH denote SLATET and SLATET’ as $\Delta$ and $\Delta^{\prime}$ respectively. When we define the SLATET, we use $D_{0}(0)=D_{0}(1)$, which follows from the sharp assignment of the instrument: $v_{10}=v_{00}$.:
$SLATET$ $=E[Y(1)-Y(0)|Z=1,T=1,D_{1}(1)>D_{0}(0)],$
$SLATET^{\prime}$ $=E[Y(1)-Y(0)|Z=0,T=1,D_{1}(0)\neq D_{0}(0)].$
This assumption requires that the treatment effects among the switchers should be the same between exposed and unexposed groups, which dCDH call the homogeneous treatment effect assumption. In Appendix E.4, we present the decomposition result for the SLATET, and show that we cannot straightforwardly interpret this assumption as requiring the homogeneous treatment effect between the two groups. In the following discussion, we therefore treat Assumptions 26-30 as dCDH’s main identifying assumptions.
E.2   Strong restrictions under Fuzzy DID designs
In this subsection, we show that dCDH’s identifying assumptions ex ante impose strong restrictions on treatment adoption behavior across units. First, we prove some lemmas implied by dCDH’s assumptions. Based on these lemmas, we then create the tables that present all the types in exposed and unexposed groups under Fuzzy DID designs. These tables show that dCDH’s assumptions ex ante exclude many types in both groups and these restrictions are heavily assigned on an unexposed group. At the end, we provide the empirical implication of the distinction between dCDH’s restrictions and ours.
First, we show that Assumption 28 implies the following assumption.
Assumption 32 (Monotone change in treatment in exposed and unexposed groups).
For all $z\in\{0,1\}$,

### v6

@@ Appendix C Decomposing the TWFEIV estimand in Oreopoulos (2006)
In this appendix, we quantify the bias terms in the TWFEIV estimand, arising from the bad comparisons in Oreopoulos (2006). Let $e=1947$ and $U=1957$ denote England and Northern Ireland respectively, and let $N_{e,t}$ denotes the sample size for England in cohort $t$. Let $R_{e,t}\equiv\frac{1}{N_{e,t}}\sum_{i}^{N_{e,t}}R_{e(i),t}$ denote the sample mean of random variable $R_{i,t}$ in region $e$ ($e\in\{1947,1957\}$) and cohort $t$. Here, $R_{e(i),t}$ indicates that the unit $i$ belongs to cohort $e$.
First, we decompose the TWFEIV estimator $\hat{\beta}_{IV}$ as follows [^23] For the detailed calculation steps, see the proof of Lemma 7 in Miyaji (2024). In our decomposition, the reference period ($t=1$) in the proof is $t=1946$.:
@@
and $\hat{Z}_{e,t}$ is the residuals from regressing $Z_{i,t}$ on the cohort and Northern Ireland fixed effects. [^24] We can write $\hat{Z}_{i,t}=\hat{Z}_{e,t}$ because $Z_{i,t}$ only varies across cohorts and regions.
Here, to clarify the interpretation of the TWFEIV estimand, we assume the parallel trends assumption in the treatment and the outcome over the entire sample period from $t=1936$ to $t=1965$. Then, by using the similar arguments in the proof of Theorem 5 for multiple time periods repeatedly, we obtain the following decomposition result for the TWFEIV estimand in Oreopoulos (2006):
$\beta_{IV}=$ $\sum\limits_{U-1\geq t\geq e}w_{e,t}^{1}\cdot CACRT_{e,t}+\sum\limits_{t\geq U}w_{e,t}^{2}\cdot\Delta_{e,t},$ (15)
@@
$\Delta_{e,t}=\displaystyle\frac{CAET^{1}_{e,t}\cdot CACRT_{e,t}-CAET^{1}_{U,t}\cdot CACRT_{U,t}}{CAET^{1}_{e,t}-CAET^{1}_{U,t}},$
@@
$=\frac{E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot(CAET^{1}_{e,t}-CAET^{1}_{U,t})}{\sum\limits_{U-1\geq t\geq e}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot CAET^{1}_{e,t}+\sum\limits_{t\geq U}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot(CAET^{1}_{e,t}-CAET^{1}_{U,t})},$ (19)
@@
where $n_{e,t}$ and $E[\hat{Z}_{i,t}|E_{i}=e]$ represent the population share and residuals for England in cohort $t$, respectively.
In Oreopoulos (2006), we can only identify $CACRT_{e,t}$ between $1947$ and $1956$ because the policy change occurred in England in $1947$ and in Northern Ireland in $1957$. This implies that each $\Delta_{e,t}$ in equation (15) is the bias term arising from the bad comparisons between $1957$ and $1965$ performed by the TWFEIV regression.
@@
Figure 3 plots the weight and the corresponding estimate for each $CACRT_{e,t}$ and $\Delta_{e,t}$ (bias term). The consistent estimators for each $w^{1}_{e,t}$, $w^{2}_{e,t}$, $CACRT_{e,t}$ and $\Delta_{e,t}$ are constructed from the sample analogue, exploiting $\widehat{WDID}_{e,t}$ and $(D_{e,t}-D_{e,1946})-(D_{U,t}-D_{U,1946})$.
Figure 3 shows that all the CACRTs are positive and positively weighted, whereas all the bias terms are positive and negatively weighted. This indicates that the TWFEIV estimand is negatively biased in the setting of Oreopoulos (2006).
@@
Figure 3: Decomposition result for the TWFEIV estimand in Oreopoulos (2006). Notes: This figure plots the estimated weights and the corresponding estimates for each $CACRT_{e,t}$ and $\Delta_{e,t}$ (bias term). The closed squares are the weights and the corresponding estimates for each $CACRT_{e,t}$. The closed triangles are the weights and the corresponding estimates for each $\Delta_{e,t}$ (bias term).
Appendix D Comparing DID-IV with Fuzzy DID
In this section, we compare the DID-IV with the Fuzzy DID proposed in de Chaisemartin and D’Haultfœuille (2018) (henceforth, “dCDH”). Although both papers consider the same empirical setting and the same Wald-DID estimand, they differ in their target parameters and associated identifying assumptions. This section clarifies the implications of these differences for treatment adoption behavior, the interpretation of the target parameter, and the use of the Wald-DID estimand.
D.1  Similarities and differences between this paper and dCDH
In this subsection, we first clarify the similarities and differences between this paper and dCDH with respect to the empirical setting, the target parameter, the estimand, and the identifying assumptions. This subsection provides the foundation for the subsequent comparison.
Empirical setting
In this paper, we mainly consider a two-period, two-group setting in which one group becomes exposed to the instrument in the second period (the exposed group), while the other group is never exposed to the instrument in either period (the unexposed group). The key features of this setting are that (i) the instrument is assigned only in the second period (sharp assignment of the instrument), and (ii) the group indicator captures exposure to the instrument. We refer to this setup as the “DID-IV” setting.
Under this definition, dCDH also consider the DID-IV setting. In particular, the group variable $G$ in dCDH plays the role of the instrument $Z$ in this paper. This follows from the fact that $G$ is included in their treatment participation equation:
$D=\mathbf{1}\{V\geq v_{GT}\}$
(see Assumption 3 in dCDH), which corresponds to condition (ii) of the DID-IV setting. Condition (i) is also assumed in dCDH: they impose $v_{10}=v_{00}$ throughout their proofs (see page 1019 of dCDH), implying that the treatment participation thresholds at time $T=0$ are identical across groups $G=0$ and $G=1$.
Here, it is useful to clarify the distinction between the empirical setting and the design in dCDH. In dCDH, the Fuzzy DID design is described as a situation in which the treatment rate in one group increases more than that in another group (see Assumption 1 in dCDH). However, this characterization corresponds to an identifying assumption, rather than a primitive feature of the empirical setting. Independently of Assumption $1$, dCDH also define the Fuzzy DID setting as $D\neq G\times T$. This definition corresponds to the DID-IV setting once combined with the restriction $v_{10}=v_{00}$.
Based on the above discussion, in what follows, we replace $G$ with $Z$ and refer to the groups $Z=1$ and $Z=0$ as the exposed and unexposed groups, respectively.
The target parameter
In this paper, we consider the LATET as our target parameter. By contrast, dCDH focus on the switcher local average treatment effect on the treated (SLATET) defined below.
Definition.
$SLATET$ $\equiv E[Y(1)-Y(0)|Z=1,T=1,D_{0}(0)<D_{1}(1)]$
$=E[Y(1)-Y(0)|Z=1,T=1,SW],$
where we use the restriction $v_{10}=v_{00}$ in dCDH, which implies $D_{0}(1)=D_{0}(0)$.
Here, to ease interpretation, we slightly modify the notation in dCDH. Specifically, we define $D_{t}(Z)=\mathbf{1}\{V\geq v_{Zt}\}$ as the treatment status at time $T=t$ for group $Z$, rather than using the notation $D(t)$ in dCDH.
This parameter measures the treatment effects for units who belong to the exposed group ($Z=1$) and switch into treatment at time $T=1$ (switchers). In subsection D.3, we clarify the differences between the SLATET and the LATET considered in this paper.
The estimand
Both papers consider the same Wald-DID estimand. However, the stance regarding the use of the Wald-DID estimand differs between this paper and dCDH. Specifically, this paper views the Wald-DID estimand as the natural estimand for identifying the LATET. By contrast, dCDH point out that the Wald-DID estimand identifies the SLATET only if the treatment effect is stable over time, and propose alternative estimands that avoid this restriction. We return to this point in subsection D.4.
The identifying assumptions
dCDH impose the following identifying assumptions for the Wald-DID estimand to identify the SLATET. For detailed discussions of each assumption, see dCDH.
@@ D.1  Similarities and differences between this paper and dCDH
$D=\mathbf{1}\{V\geq v_{ZT}\},\hskip 2.84526pt\text{with}\hskip 5.69054ptV\mathop{\perp\!\!\!\!\perp}T|Z.$
@@
Again, in Assumption 35, we adopt the notation $D_{t}(Z)=\mathbf{1}\{V\geq v_{Zt}\}$, rather than $D(t)$ used in dCDH, to facilitate interpretation.
We now state the detailed connections between the identifying assumptions in dCDH and those in this paper. Because dCDH consider a repeated cross-section setting, when comparing the identifying assumptions across the two papers, we refer to Assumptions 19-24.
First, both dCDH and this paper impose condition (21) in Assumption 31, which corresponds to the standard relevance condition. Next, both dCDH and this paper impose Assumption 20. Specifically, Assumption 33 (treatment participation equation) in dCDH implies Assumption 20 in this paper (Vytlacil (2002)).
Moreover, dCDH implicitly impose Assumptions 19 and 21 when formalizing potential outcomes and treatment participation. Specifically, dCDH do not introduce instrument paths in their framework and define potential outcomes as $Y(D)$, which implies Assumption 19. In addition, as shown in Assumption 33, dCDH assume that the treatment participation equation depends only on the current instrument status. Combined with sharp assignment of the instrument ($v_{10}=v_{00}$), this implies Assumption 21.
The differences between dCDH’s assumptions and ours are as follows. First, dCDH impose the parallel trends assumption in untreated outcomes (Assumption 34), whereas we impose the parallel trends assumption in unexposed outcomes. Second, dCDH assume stable treatment rates in the unexposed group (Assumption 32), whereas we assume the parallel trends in treatment between exposed and unexposed groups. Third, dCDH impose an independence assumption between unobserved heterogeneity $V$ and time $T$ conditional on $Z$ in Assumption 33, whereas we do not impose such a restriction. Fourth, the treatment participation equation in Assumption 33 is stronger than Assumption 20 in this paper. In particular, Assumption 33 imposes the monotonicity with respect to time $T$, in addition to the monotonicity with respect to instrument $Z$ (see Lemma 1 in Subsection D.2). Finally, dCDH impose condition (20) in Assumption 31, which requires that the treatment rate in the exposed group increases between the two periods, whereas we do not impose this condition.
In the next subsection, we examine how these differences imply different restrictions on treatment adoption behavior across the two designs.
Remark 3.
@@
where we define252525dCDH denote SLATET and SLATET’ as $\Delta$ and $\Delta^{\prime}$ respectively. When we define the SLATET, we use $D_{0}(0)=D_{0}(1)$, which follows from the sharp assignment of the instrument: $v_{10}=v_{00}$.:
$SLATET$ $=E[Y(1)-Y(0)|Z=1,T=1,D_{0}(0)<D_{1}(1)],$
$SLATET^{\prime}$ $=E[Y(1)-Y(0)|Z=0,T=1,D_{0}(0)\neq D_{1}(0)].$
This assumption requires that the treatment effects among the switchers should be the same between exposed and unexposed groups, which dCDH call the homogeneous treatment effect assumption. In Appendix D.3, we present the decomposition result for the SLATET, and show that we cannot straightforwardly interpret this assumption as requiring the homogeneous treatment effect between the two groups. In the following discussion, we therefore treat Assumptions 31-35 as dCDH’s main identifying assumptions.
D.2  Restrictions on treatment adoption behavior
In this subsection, we compare the restrictions on treatment adoption behavior between the two designs. Since both this paper and dCDH impose Assumption 21 (see the previous section), we adopt the notation $(D_{0}(0),D_{1}(0),D_{1}(1))$ for potential treatment choices in the following. Note that while we do not use $D_{0}(1)$ in DID-IV designs due to the sharp assignment of the instrument, we can define $D_{0}(1)=\mathbf{1}\{V\geq v_{10}\}$ in dCDH’s notation. However, since dCDH assume $v_{10}=v_{00}$, we have $D_{0}(1)=D_{0}(0)$ almost surely. We use this fact repeatedly in the following discussion.
D.2.1 Preparation
We first describe the heterogeneity in treatment adoption behavior under the DID-IV setting. This allows us to clarify which latent adoption types are ruled out by the identifying assumptions in Fuzzy DID, and how these restrictions differ from those imposed in DID-IV. To do so, we introduce the following notation. Let $G^{T}=(D_{0}(0),D_{1}(0))$ denote the group variable that represents the treatment adoption process across the two periods when the instrument is $Z=0$ in the second period. That is, $G^{T}$ captures whether an individual would start, stop, or never adopt treatment over time in the absence of the instrument. Analogous to $G^{Z}=(D_{1}(0),D_{1}(1))$ (which was introduced in Section 2.1 for the panel data case), we define $G^{T}=(0,0)\equiv NT^{T}$ as time never-takers, $G^{T}=(1,1)\equiv AT^{T}$ as time always-takers, $G^{T}=(0,1)\equiv CM^{T}$ as time compliers, and $G^{T}=(1,0)\equiv DF^{T}$ as time defiers.
Using the group variables $G^{Z}$ and $G^{T}$, we can partition units into eight types within each group. Specifically, we obtain the following tables for the exposed and unexposed groups, respectively.
Table 7: Exposed group ($z=1$)
observed counterfactual